ROTATIONAL ANALYSIS OF THE C̃^{2}A_{1} - X̃^{2}A_{1} TRANSITION OF SrNH_{2}

\underline{P. M. SHERIDAN}, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; M. J. DICK, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; J. G. WANG and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

A high-resolution spectrum of the C̃^{2}A_{1} - X̃^{2}A_{1} transition of SrNH_{2} has been recorded using a jet cooled/laser ablation spectrometer. SrNH_{2} was created by the reaction of UV-ablated Sr atoms with a 15% mixture of NH_{3} in argon and detected using laser excitation spectroscopy. Low J transitions arising from both the K′_{a} = 0 ← K′′_{a} = 0 and K′_{a} = 1 ← K′′_{a} = 1 sub-bands have been observed. Rotational and fine structure parameters have been determined for both states. The spin-rotation constants for the C̃^{2}A_{1} state have been calculated using the pure precession model and will be compared to those determined in the fit. In addition, these spectroscopic constants will be compared to those of the previously measured Ã^{2}B_{1} and B̃^{2}B_{2} states of SrNH_{2}. Structural parameters for all states will also be discussed.